How many kg of a 45% copper alloy must be added to 62kg of a 60% copper alloy to form an alloy, which is 50% copper.
Answer by stanbon(75887)

How many kg of a 45% copper alloy must be added to 62kg of a 60% copper alloy to form an alloy, which is 50% copper.
======================
Equation:
copper + copper = copper
0.45x + 0.60*62 = 0.50(x+62)
------
45x + 60*62 = 50x + 50*62
-------
5x = 10*62
x = 124 kg (amt. of 45% alloy needed)
--------------
Cheers,
Stan H.
---------------